Simsons inc. can sell 2800 pairs of boots per week if they charge $140 sales drop to 2750 pairs per week if they charge $150. Write a formula for demand function p=D(q)


1
Expert's answer
2021-02-09T07:02:17-0500

The demand function can be written as follows:


"p=mQ+b."

Let's first find the slope of the demand function:


"m=\dfrac{y_2-y_1}{x_2-x_1}=\dfrac{150-140}{2750-2800}=-\dfrac{10}{50}=-0.2."

Then, we can write the equation of the demand function:


"(y-y_1)=m(x-x_1),""p-140=-0.2(Q-2800),""p-140=-0.2Q+560,""p=-0.2Q+700."